# thanks: ISTI Fellow

### v1

@@ I Introduction
The detection of gravitational waves (GWs) emitted by binary neutron star (BNS) merger GW170817 [1, 2] and, subsequently, electromagnetic (EM) radiation from its associated counterpart AT2017gfo [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] introduced exciting prospects for studying the universe using multimessenger astronomy. The GWs emitted by BNS mergers carry information about the pre-merger binary, such as the masses, spins, and tidal deformabilities of the merging neutron stars. Likewise, the EM radiation from the counterpart, or “kilonova” [20, 21], conveys details about the material that was ejected during and after the merger [22], tying directly to the fate of the merger remnant [23, 24, 25], the dense matter equation of state (EOS) [26, 27], and r-process nucleosynthesis [28, 29, 30].
@@
Following the GW170817/AT2017gfo detection, recent studies have performed numerous numerical relativity simulations of binary neutron star mergers (e.g. [31, 32, 33, 34, 35, 36, 37, 38, 39, 33, 40, 34, 41, 32, 42, 43, 44]). Using these simulations, certain studies identified fitting formulae that connect ejecta properties to the parameters of the binary that are easily extractable from the GW signal [45, 46, 47]. These fits have been used in previous multimessenger analyses to place stronger constraints on unknown quantities like the EOS [46, 48, 49], particularly in the case of GW190425, a black-hole-neutron-star merger whose ejected mass had to be inferred using one of these fits due to the lack of an electromagnetic counterpart [50].
@@
Previous work has identified that these fits to numerical relativity simulations exhibit variable validity in different regions of binary parameter space [51]. Moreover, the underlying simulations used to build these fits only cover a very limited range of possible binary merger configurations. Nonetheless, as we will demonstrate in this work, large-scale multimessenger inference calculations must employ these fits far outside the regions in which they have been calibrated. The need for using these fits outside of their safely-calibrated region has not been emphasized in the many previous studies that have employed them (e.g. [49]).
@@
Motivated by a recently developed surrogate model for kilonova emission that provides the most constraining ejecta information to date [52], in conjunction with previous studies examining the agreement between ejecta fits (e.g. [51]), in this study we examine how these ejecta fits inform multimessenger inference of the binary parameters. We emphasize that our kilonova model is by no means complete in the sense that many systematic uncertainties are still omitted. Recent studies have identified multiple additional sources of systematic uncertainty along the merger-to-kilonova modeling chain [53, 54, 55]. Nevertheless, our surrogate model’s ability to tightly constrain ejecta parameters enables us to investigate the systematic uncertainties that propagate into Bayesian inference analyses that utilize these fits. Our intent with this study is to highlight the systematic uncertainties that are propagated when using these fits for more informed use in future studies.
@@ II Methodology
II.1 Ejecta fits to numerical relativity simulations
@@
We employ the forward model ejecta fits from Refs. [45, 46, 47], hereafter referred to as KruFo20, DiCo20, and Nedora21, respectively. These fits to numerical relativity simulations estimate the ejecta properties, namely dynamical and wind ejecta mass and velocity, given binary properties (mass ratio $q$ and neutron star radius $R_{\rm{1.4}}$) as inputs. In summary, KruFo20 present fitting formulae based on 52 numerical relativity simulations of binary neutron star mergers compiled from Refs. [33] and [34]. DiCo20 base their formulae on 73 numerical relativity simulations, using the two aforementioned sources in addition to simulations from Refs. [32] and [31]. Finally, Nedora21 employs the largest numerical relativity dataset to date for their fits, using 324 models from Refs. [35, 36, 37, 38, 39, 33, 40, 34, 41, 32, 42, 43, 44]. The implementation of these ejecta fits follows the approach outlined in Section II.F of Ref. [56], with the exception of the modifications described in this section.
@@
Figure 1: Solid, dashed, and dotted line contours corresponding to the KruFo20 dynamical ejecta mass $m_{\rm{ej}}$, disk mass $m_{\rm{disk}}$, and ejecta velocity $v_{\rm{ej}}$, respectively. Contours were calculated as a function of mass ratio $\delta=(m_{1}-m_{2})/(m_{1}+m_{2})$ and neutron star radius $R_{\rm{1.4}}$. Mass and velocity units are in $M_{\odot}$ and $c$, respectively.
@@
Each ejecta fit predicts the ejecta mass, ejecta velocity, and accretion disk mass as a function of the binary mass ratio $q$ and the neutron star radius $R_{\rm{1.4}}$. In none of the ejecta fit models is there a prescription for tying the disk mass to post-merger wind mass; instead, we map the fits’ ejecta mass prediction to the dynamical ejecta, and assume some fraction $f_{\rm{disk}}$ of the disk mass is ejecta as wind ejecta. In Figure 1, we display naive contours corresponding to representative values of these ejecta parameters as a function of the binary parameters, fixing the binary source-frame chirp mass to the observed value for GW170817. The solid, dashed, and dotted lines indicate $m_{\rm{ej}}(\delta,R_{\rm{1.4}})$, $m_{\rm{disk}}(\delta,R_{\rm{1.4}})$, and $v_{\rm{ej}}(\delta,R_{\rm{1.4}})$, respectively. The solid dynamical ejecta contours constrain the mass ratio $q$ while allowing for a broad range of radii. Conversely, the dashed disk mass contours constrain $R_{\rm{1.4}}$ to favor low values for realistic disk masses with a slight constraint on $\delta$. Critically, however, the dotted $v_{\rm{ej}}$ contours – shown at levels corresponding to our previously-inferred estimate for this quantity – exclude the conventional interpretation for GW170817, which consists of $\delta\simeq 0$ (comparable masses) and $R_{1.4}\simeq 12{\rm km}$. Instead, the $v_{\rm{ej}}$ contours follow a trendline that permits either comparable mass but large radius or conventional radius but highly asymmetric masses, both scenarios that could produce sufficiently low dynamical ejecta velocities in binary neutron star mergers.
@@
Rather than estimating the uncertainties on the ejected mass from each of our two components, we assign free parameters $\alpha_{\rm{dyn}}$ and $f_{\rm{disk}}$ to quantify these systematic uncertainties during inference. The $\alpha_{\rm{dyn}}$ parameter serves as a scale factor for the fits’ dynamical ejecta mass prediction, while $f_{\rm{disk}}$ dictates what fraction of the accretion disk is expelled as wind ejecta. Given the behavior of the dotted $v_{\rm{ej}}$ contours in Figure 1, we also include a systematic uncertainty parameter for the ejecta velocity, $\beta_{\rm\phi}$. To uniformly explore the systematic uncertainty in velocity space, we apply the reparameterization $\phi=\ln((c/v)^{2}-1)$. Our uncertainty is characterized by a uniformly distributed random variable $\beta_{\phi}$ added to $\phi$.
@@
The ejecta fits predict a total ejecta velocity $v_{ej}$ by adding in quadrature the velocities describing the ejecta within and perpendicular to the orbital plane, without explicitly declaring to which of the dynamical and wind components that ejecta belongs (see Eqs. (5)–(9) in Ref. [57]). In this study, following previous conventions (e.g. [58]), we set the dynamical ejecta velocity to the value calculated only for the perpendicular velocity component such that $v_{d}=v_{z}$ and fix the wind velocity $v_{w}=0.10c$ per the results of our previous analysis in Ref. [52]. This choice is similar to previous approaches, e.g. as in Ref. [56], where the wind velocity $v_{w}$ was set to $0.08c$ as informed by Ref. [33].
@@
The employed fits also have a lower limit on the accretion disk mass of $\log M_{\rm{disk}}\geq-3$, but no defined upper limit. As certain combinations explored during sampling may yield binary parameters outside of the realm of validity for the ejecta fits, we set an associated restriction on the wind mass such that $\log m_{w}\leq-1$.
II.2 Kilonova simulations and surrogate model
We assume a two-component kilonova model consisting of lanthanide-rich, equatorial dynamical ejecta and lanthanide-poor, axial wind ejecta as described in [59, 60] and motivated by numerical simulations [22, 61]. Each component is assumed to be homologously expanding and parameterized by a mass and velocity such that $M_{\rm{d}}$, $v_{\rm{d}}$ and $M_{\rm{w}}$, $v_{\rm{w}}$ describe the dynamical and wind components’ masses and averaged velocities, respectively. The morphology for the dynamical component is an equatorially centered torus, while the wind component is represented by an axially-centered peanut component; Figure 1 of [59] displays the torus-peanut, or “TP,” schematic corresponding to the morphologies employed in this work; see Ref. [60] for a detailed definition.
The lanthanide-rich dynamical ejecta stems from $r$-process nucleosynthesis in the neutron-rich ejected material, characterized by a low electron fraction ($Y_{\rm{e}}\equiv n_{\rm{p}}/(n_{\rm{p}}+n_{\rm{n}})$) of $Y_{e}=0.04$, with elements reaching the third $r$-process peak ($A\sim 195$). The wind ejecta originates from higher $Y_{e}=0.27$ that encapsulates elements between the first ($A\sim 80$) and second ($A\sim 130$) $r$-process peaks. The detailed breakdown of the elements in each component can be found in Table 2 of Ref. [59].
The SuperNu outputs are observing-angle-dependent, simulated spectra $F_{\lambda,\rm sim}$, post-processed to a source distance of $10$ pc, in units of erg s-1 cm-2 Å-1. The spectra are binned into 1024 equally log-spaced wavelength bins spanning $0.1\leq\lambda\leq 12.8$ microns. For the purposes of this work, we consider the light curves for the 2MASS $grizy$ and Rubin Observatory $JHK$ broadband filters. As we only consider anisotropic simulations in this study, unless otherwise noted, we extract simulated light curves using 54 angular bins, uniformly spaced in $\cos\theta$ over the range $-1\leq\cos\theta\leq 1$. The angle $\theta$ is taken between the line of sight and the symmetry axis as defined in Equation (2) of Ref. [52], with further description of the treatment of observing angle found therein.
We interpolate over the library of SuperNu simulations described in [72, 52] using the multi-layer perceptron (MLP) described in [52]. In summary, the MLP was trained on $\sim 250$ light-curve simulations evaluated at $264$ log-spaced times between $0.125$ and $37.24$ days for $54$ viewing angles equally spaced in $\cos\theta$ for $\theta$ ranging from 00 to $180^{\circ}$. We do not perform any normalization of our inputs or outputs, with ejecta parameters ranging from $-3\leq\log m/M_{\odot}\leq-1$ and $0.05\leq v/c\leq 0.3$ and light curves ranging from -18 to 8 AB magnitudes. We train a separate MLP for each of our broadband filters, training each MLP for 1000 epochs with a batch size of 32 using the Adam optimizer. Our initial learning rate is $2~{}\times~{}10^{-4}$ with a decay rate of 5% every 10 epochs.

### v2

@@ I Introduction
The detection of gravitational waves (GWs) emitted by the binary neutron star (BNS) merger GW170817 [1, 2] and, subsequently, electromagnetic (EM) radiation from its associated optical/infrared counterpart AT2017gfo [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] introduced exciting prospects for studying the universe using multimessenger astronomy. The GWs emitted by BNS mergers carry information about the pre-merger binary, such as the masses, spins, and tidal deformabilities of the merging neutron stars. Likewise, the EM radiation from the optical/infrared counterpart, or “kilonova” [20, 21], conveys details about the material that was ejected during and after the merger [22], tying directly to the fate of the merger remnant [23, 24, 25], the dense matter equation of state (EOS) [26, 27], and rapid neutron capture process (r-process) nucleosynthesis [28, 29, 30].
@@
Following the GW170817/AT2017gfo detection, recent studies have performed numerous numerical relativity simulations of BNS mergers [31, 32, 33, 34, 35, 36, 37, 38, 39, 33, 40, 34, 41, 32, 42, 43, 44]. Using these simulations, certain studies identified fitting formulae that connect BNS merger ejecta properties to the parameters of the binary that are easily extractable from the GW signal [45, 46, 47]. Mapping these ejecta properties to those relevant to kilonova model ejecta is a crucial ingredient in enabling multimessenger inferences using both GW and EM information. These fits and ejecta mappings have been used in previous multimessenger analyses to place stronger constraints on unknown quantities like the EOS [46, 48, 49], particularly in the case of GW190425, a BNS/black-hole-neutron-star merger whose ejected mass had to be inferred using one of these fits due to the lack of an electromagnetic counterpart [50]. It is important to note that the BNS merger ejecta to kilonova ejecta mapping is highly non-trivial, and an accurate one-to-one mapping still eludes the field. Numerical relativity models of BNS mergers are still computationally expensive, with recent record-breaking simulations only going out to $\sim 1.5$ s post-merger [51]. The intrinsic variability in the BNS parameter space, in addition to kilonova time and length scales being many orders of magnitude larger than the longest numerical BNS models, necessitates extrapolation or projection in the mapping of BNS merger ejecta to kilonova model ejecta.
@@
Previous work in the literature has identified that these fits to numerical relativity simulations exhibit variable validity in different regions of binary parameter space [52, 53]. Moreover, the underlying simulations used to build these fits only cover a very limited range of possible BNS merger configurations. Nonetheless, as we will demonstrate in this work, large-scale multimessenger inference calculations must employ these fits far outside the regions in which they have been calibrated to recover reasonable binary parameters.
The need for using these fits outside of their safely-calibrated region has not been emphasized in the many previous studies or frameworks that have employed them. Several (though not all) previous multimessenger inference studies [54, 55, 46, 49] present joint EOS inference with systematically incomplete ejecta fit models, by marginalizing over the ejecta velocity for all kilonova ejecta components, leaving ejecta velocity inference as a free parameter, unconstrained by any modeling. In particular, these studies neglect to verify whether the velocities required by the kilonova models to reproduce observed light curves are even qualitatively consistent with the velocity expectations from the ejecta fits. Furthermore, these analyses were carried out by reweighting GW samples during joint inference, thus carrying a convenience prior that hides the effects that arise when performing simultaneous joint inference as explored in this study.
@@
Motivated by a recently developed surrogate model for kilonova emission that provides the most constraining ejecta posteriors to date [56] (which do not necessarily encapsulate the full breadth of kilonova modeling systematics), in conjunction with previous studies examining the agreement between ejecta fits (e.g. [53]), in this study we examine how these ejecta fits inform multimessenger inference of the binary parameters. In this work, we assume a specific mapping between these ejecta types given its prior use in previous works and focus solely on the possible impact of the fitting formulae. We caution readers that our assumption of ejecta mapping is by no means complete or one-to-one, but is a convention readily assumed in the field. We emphasize that our kilonova model is also by no means complete in the sense that many systematic uncertainties are still omitted. Recent studies have identified multiple additional sources of systematic uncertainty along the merger-to-kilonova modeling chain [57, 58, 59]. Nevertheless, our surrogate model’s ability to tightly constrain ejecta parameters enables us to investigate the systematic uncertainties that propagate into Bayesian inference analyses that utilize these fits. Our approach contrasts previous studies which assume kilonova systematic uncertainties anywhere between 1 magnitude [60, 61, 54] and 5 magnitudes [62]; whereas the ejecta fit systematics have historically been washed out by the kilonova systematics, we take our surrogate model’s constrained posteriors for granted and directly examine the systematics associated primarily with the ejecta fits. Our intent with this study is to highlight the systematic uncertainties that are propagated when using these fits for more informed use in future studies.
@@ II Methodology
II.1 Kilonova simulations and surrogate model
We assume a two-component kilonova model consisting of lanthanide-rich, equatorial dynamical ejecta and lanthanide-poor, axial wind ejecta as described in [63, 64] and motivated by numerical simulations [22, 65]. Each component is assumed to be homologously expanding and parameterized by a mass and velocity such that ($m_{\rm{d}}$, $v_{\rm{d}}$) and ($m_{\rm{w}}$, $v_{\rm{w}}$) describe the dynamical and wind components’ masses and averaged velocities, respectively. The morphology for the dynamical component is an equatorially centered torus, while the wind component is represented by an axially-centered peanut component; Figure 1 of [63] displays the torus-peanut, or “TP,” schematic corresponding to the morphologies employed in this work; see Ref. [64] for a detailed definition.
The lanthanide-rich dynamical ejecta stems from $r$-process nucleosynthesis in the neutron-rich ejected material, characterized by a low electron fraction ($Y_{\rm{e}}\equiv n_{\rm{p}}/(n_{\rm{p}}+n_{\rm{n}})$) of $Y_{e}=0.04$, with elements reaching the third $r$-process peak ($A\sim 195$). Our dynamical ejecta explores the nucleosynthesis resulting from the full $r$-process; however, recent studies have shown that the dynamical ejecta can be less neutron rich when considering neutrino interactions (see, e.g., Ref. [66]). The wind ejecta originates from higher $Y_{e}=0.27$ that encapsulates elements between the first ($A\sim 80$) and second ($A\sim 130$) $r$-process peaks. The detailed breakdown of the elements in each component can be found in Table 2 of Ref. [63].
The SuperNu outputs are observing-angle-dependent, simulated spectra $F_{\lambda,\rm sim}$, post-processed to a source distance of $10$ pc, in units of erg s-1 cm-2 Å-1. The spectra are binned into 1024 equally log-spaced wavelength bins spanning $0.1\leq\lambda\leq 12.8$ microns. For the purposes of this work, we consider the light curves for the Rubin Observatory grizygrizy and 2MASS $JHK$ broadband filters. As we only consider anisotropic simulations in this study, unless otherwise noted, we extract simulated light curves using 54 angular bins, uniformly spaced in $\cos\theta$ over the range $1\leq\cos\theta\leq 1$. The angle $\theta$ is taken between the line of sight and the symmetry axis ($z=0$) as defined in Equation (2) of Ref. [56], with further description of the treatment of observing angle found therein.
We interpolate over the library of SuperNu simulations described in [77, 56] using the multi-layer perceptron (MLP) described in [56]. In summary, the MLP was trained on $\sim 250$ light-curve simulations evaluated at $264$ log-spaced times between $0.125$ and $37.24$ days for $54$ viewing angles equally spaced in $\cos\theta$ for $\theta$ ranging from 0 to $180^{\circ}$. We do not perform any normalization of our inputs or outputs, with ejecta parameters ranging from $3\leq\log m/M_{\odot}\leq-1$ and $0.05\leq v/c\leq 0.3$ and light curves ranging from -18 to 8 AB magnitudes. We train a separate MLP for each of our broadband filters, training each MLP for 1000 epochs with a batch size of 32 using the Adam optimizer. Our initial learning rate is $2~\times~10^{-4}$ with a decay rate of 5% every 10 epochs.
II.2 Ejecta fits to numerical relativity simulations
@@
We employ the forward model ejecta fits from Refs. [45, 46, 47], hereafter referred to as KruFo20, DiCo20, and Nedora21, respectively. These fits to numerical relativity simulations estimate the ejecta properties, namely dynamical and wind ejecta mass and velocity, given binary properties (mass ratio $q$ and neutron star radius $R_{\rm{1.4}}$) as inputs. In summary, KruFo20 present fitting formulae based on 52 numerical relativity simulations of binary neutron star mergers compiled from Refs. [33] and [34]. DiCo20 base their formulae on 73 numerical relativity simulations, using the two aforementioned sources in addition to simulations from Refs. [32] and [31]. Finally, Nedora21 employs the largest numerical relativity dataset to date for their fits, using 324 models from Refs. [35, 36, 37, 38, 39, 33, 40, 34, 41, 32, 42, 43, 44]. The implementation of these ejecta fits follows the approach outlined in Section II.F of Ref. [78], with the exception of the modifications described in this section. These fits neglect the effects of NS spin on the amount of ejected material, which can impact the amount of ejecta by factors of order unity for rapidly spinning NS $(\chi\gtrsim 0.1)$; see, e.g., [32, 79, 80].
@@
Figure 1: Solid, dashed, and dotted line contours corresponding to the KruFo20 dynamical ejecta mass $m_{\rm{ej}}$, disk mass $m_{\rm{disk}}$, and ejecta velocity $v_{\rm{ej}}$, respectively. Contours were calculated as a function of mass asymmetry $\delta=(m_{1}-m_{2})/(m_{1}+m_{2})$ and neutron star radius $R_{\rm{1.4}}$. Mass and velocity units are in $M_{\odot}$ and $c$, respectively.
@@
Each ejecta fit predicts the ejecta mass, total ejecta velocity, and accretion disk mass as a function of the binary mass ratio $q$ and the neutron star radius $R_{\rm{1.4}}$. We relate $R_{\rm{1.4}}$ to $\tilde{\Lambda}$ as $R_{\rm{1.4}}\simeq 13.4(\tilde{\Lambda}/800)^{1/6}$ km per Eq. (18) in Ref. [81]. In none of the ejecta fit models is there a prescription for tying the disk mass to post-merger wind mass; instead, we map the fits’ ejecta mass prediction to the dynamical ejecta, and assume some fraction $f_{\rm{disk}}$ of the disk mass is ejected as wind ejecta. In Figure 1, we display naive contours corresponding to representative values of these ejecta parameters as a function of the binary parameters, fixing the binary source-frame chirp mass to the observed value for GW170817. The solid, dashed, and dotted lines indicate $m_{\rm{ej}}(\delta,R_{\rm{1.4}})$, $m_{\rm{disk}}(\delta,R_{\rm{1.4}})$, and $v_{\rm{ej}}(\delta,R_{\rm{1.4}})$, respectively, where $\delta=(m_{1}-m_{2})/(m_{1}+m_{2})=(1-q)/(1+q)$ indicates the mass asymmetry in the binary. For reference, for GW170817-like systems as the one shown in Figure 1, $\delta\sim 0.1\implies q\sim 0.81$ while $\delta\sim 0.2\implies q\sim 0.66$. The $\delta$ space explored in our study allows for a substantial fraction of extremely asymmetric configurations, which we will show are necessarily explored by the ejecta fits to reconcile EM and GW observations (assuming the constraints presented by our kilonova model) without the use of additional astrophysical priors. The solid dynamical ejecta contours constrain the mass ratio $q$ while allowing for a broad range of radii. Conversely, the dashed disk mass contours constrain $R_{\rm{1.4}}$ to favor low values for realistic disk masses ($m_{\rm{disk}}/M_{\odot}<0.45$, as in, e.g., Ref. [82]) with a slight constraint on $\delta$. Critically, however, the dotted $v_{\rm{ej}}$ contours – shown at levels corresponding to our previously-inferred estimate (0.100.20c0.10-0.20c, [77, 83]) for this quantity – exclude the conventional interpretation for GW170817, which consists of $\delta\simeq 0$ (comparable masses) and $R_{1.4}\simeq 12{\rm km}$. Instead, the $v_{\rm{ej}}$ contours follow a trendline that permits either comparable mass but large radius or conventional radius but highly asymmetric masses, both scenarios that could produce sufficiently low dynamical ejecta velocities in binary neutron star mergers.
@@
Rather than estimating the uncertainties on the ejected mass from each of our two components, we assign free parameters $\alpha_{\rm{dyn}}$ and $f_{\rm{disk}}$ to quantify these systematic uncertainties during inference. The $\alpha_{\rm{dyn}}$ parameter serves as a scale factor for the fits’ dynamical ejecta mass prediction, while $f_{\rm{disk}}$ dictates what fraction of the accretion disk is expelled as wind ejecta. Given the behavior of the dotted $v_{\rm{ej}}$ contours in Figure 1, we also introduce a systematic uncertainty parameter for the ejecta velocity, $\beta_{\rm\phi}$. To uniformly explore the systematic uncertainty in velocity space, we apply the reparameterization $\phi=\ln((c/v)^{2}-1)$. Our uncertainty is characterized by a uniformly distributed random variable $\beta_{\phi}$ added to $\phi$.
@@
The ejecta fits predict a total ejecta velocity $v_{ej}$ by adding in quadrature the velocities describing the ejecta within and perpendicular to the orbital plane, without a direct one-to-one mapping to the standard kilonova dynamical and wind ejecta components (see Eqs. (5)–(9) in Ref. [60]). In this study, to ensure maximal consistency with established methodology in the literature, we inherit previous conventions. As in Ref. [54], we set the dynamical ejecta velocity to that of the total ejecta velocity such that $v_{d}=v_{ej}$ and fix the wind velocity $v_{w}=0.10c$ per the results of our previous analysis in Ref. [56]. The choice to fix $v_{w}$ is similar to previous approaches, e.g. as in Ref. [78], where the wind velocity $v_{w}$ was set to $0.08c$ as informed by Ref. [33]. We caution readers that in attributing $v_{ej}=v_{d}$ as a makeshift connection between binary and ejecta parameters, we stray away from the expected evolution of key quantities (e.g. matter density, temperature) as a function of their total expansion velocity.
@@
The employed fits also have a lower limit on the accretion disk mass of $\log M_{\rm{disk}}\geq-3$, but no defined upper limit. As certain combinations explored during sampling may yield ejecta parameters outside of the realm of validity of our surrogate model, we set an associated restriction on the wind mass such that $\log m_{w}\leq-1$. The same upper limit applies to the dynamical ejecta to ensure ejecta samples stay within the region of validity.